OpenStudy (bookworm14):

Please Help, I am so confused! I will post it below in the comments, thank you so much for taking the time to look at it!!

OpenStudy (bookworm14):

Help Please?

OpenStudy (babyx3boo):

The first formula is for the sphere one and the second one for the cone. Just plug the numbers into the formula.

OpenStudy (rea201):

ok so because they have the same volume you need to set the two equations to equal each other and subtitue for the radius then solve for h

OpenStudy (bookworm14):

I know which formula is which, I am just not sure what to do with it (or where to plug the numbers in at)

OpenStudy (babyx3boo):

in place of the radius (r)

OpenStudy (rea201):

so \[\frac{ 4 }{ 3 }\times \pi \times (2a)^3=\frac{ 1 }{ 3 }\times \pi \times (2a)^2\]

OpenStudy (bookworm14):

I plugged it and got 36a(pi) and 3/4a(pi), divided and 26 divided by 3/4 and got 27, am i on the right track?

OpenStudy (bookworm14):

*36 not 26 typo

OpenStudy (babyx3boo):

oh dang i tried to do the problem, yea i see your confusion let me see if i can find a way myself .

OpenStudy (bookworm14):

Ok, im working on other problems, while we are doing this one, so don't rush just take your time :) lol

OpenStudy (babyx3boo):

I got it!

OpenStudy (bookworm14):

Hallelujah! :D lol

OpenStudy (rea201):

\[4/3 \times \pi \times (3a)^3=1/3 \times \pi \times (2a)^2 \rightarrow 36a^3*\pi=4/3 \times \pi \times a^2\]

OpenStudy (rea201):

\[27a=h\]

OpenStudy (babyx3boo):

lol ok i don't know to explain it ; but let's start with the sphere one; formula: V=4/3(3.14)(r)^3 meaning V= 4/3(3.14)(3a^3) do that and see what you get.

OpenStudy (babyx3boo):

i mean (3a)^3 sorry lol

OpenStudy (bookworm14):

I got: including pi (3.14) = 113.04 excluding pi (31.4)= 36

OpenStudy (babyx3boo):

Okay include the pi always and 113.04 meaning V= 113.04 a^3 don't forget the a^3 . Since you know that the volumes are the same for each formula you're going to go head and plug in the numbers in the cone formula. Cone: V= 1/3(3.14)(r)^2h 113.04a^3= 1/3(3.14)(2a)^2h multiply the numbers to the right side and tell me what you get.

OpenStudy (babyx3boo):

including pi. and the letters.

OpenStudy (bookworm14):

4.186(a)(h) ?

OpenStudy (babyx3boo):

remember the a is square but yes; but if you notice in the calculator (if you are using one) the 6 goes on right? well we will go back to that . so you're trying to find H. and you got 113.04a^3= 4.1866666666666667a^2h ; you're trying to get h by itself so divide both sides by 4.1866666666667a^2 so you can get h by itself.

OpenStudy (babyx3boo):

tell me what you get?

OpenStudy (bookworm14):

27a^2

OpenStudy (babyx3boo):

you're close but no

OpenStudy (babyx3boo):

are you using a calculator?

OpenStudy (bookworm14):

yes, but I changed my mind I think it is 27a not 27a^2 because if you divide then you subtract exponents... at least I think that is what this means? (correct me if im wrong plz lol)

OpenStudy (babyx3boo):

Yessss! there ya go :)

OpenStudy (bookworm14):

Oh goodness, I knew the whole exponent thing my brain was just slowing down on me all of a sudden haha, thank you so much for helping me work all the way through!!! Now I know how to defeat....the beast! >:D lol

OpenStudy (babyx3boo):

lol no problem :) if anything let me know

OpenStudy (bookworm14):

Ok thanks!